# On Mixed-Precision Iterative Methods and Analysis for Nearly Completely Decomposable Markov Processes

### v1

@@ Abstract
In this paper we consider the problem of computing the stationary distribution of nearly completely decomposable Markov processes, a well-established area in the classical theory of Markov processes with broad applications in the design, modeling, analysis and optimization of computer systems. We design general classes of algorithmic solution approaches that exploit forms of mixed-precision computation to significantly reduce computation times and that exploit forms of iterative approximate methods to mitigate the impact of inaccurate computations, further reduce computation times, and ensure convergence. Then we derive a mathematical analysis of our general algorithmic approaches that establishes theoretical results on approximation errors, convergence behaviors, and other algorithmic properties. Numerical experiments demonstrate that our general algorithmic approaches provide significant improvements in computation times over the most efficient existing numerical methods.
@@ 1 Introduction
Markov processes often play an important role in the design and mathematical modeling, analysis and optimization of a wide range of computer systems. A critically important goal in the mathematical modeling, analysis and optimization of such computer systems and their dynamic behavior concerns the need to determine the stationary distribution of the corresponding Markov process, from which various measures and quantities of interest can be directly obtained to facilitate and support the design, analysis and optimization of computer systems and applications.
More precisely, consider a discrete-time Markov process $\{X(s)\,;\,s\in\mathbb{Z}_{+}\}$ defined on the state space $[n]:=\{1,\ldots,n\}$, and let ${\mathbf{P}}\in\mathbb{R}^{n\times n}$ and $\bm{\pi}:=(\pi_{1},\ldots,\pi_{n})$ respectively denote its transition probability matrix and stationary distribution. Here, $\mathbb{Z}_{+}$ and $\mathbb{R}$ denote the set of nonnegative integers and real numbers, respectively. Assuming this Markov process is irreducible and ergodic, then the invariant probability vector $\bm{\pi}$ exists and is uniquely determined as the solution of [25, 26]
@@
One particularly important class of structural properties often arising in such Markov processes concerns the so-called “aggregation of variables” [22] in which all the states (variables) of the system can be grouped into a relatively small number of relatively large clusters such that the magnitude of transitions between states within a cluster overwhelmingly dominate the relatively tiny magnitude of transitions between states in different clusters. Markov processes with these structural properties, called nearly completely decomposable (NCD), make it possible to effectively investigate the system and its dynamics by exploiting the analysis of interactions among the states within each individual cluster without reference to the interactions among the states across clusters and by exploiting the analysis of interactions among the states across different clusters without reference to the interactions among the states within clusters. Such NCD properties arise naturally in many computer systems and applications due to the prevalence of levels of modular abstraction or hierarchies of components in the design of these computer systems and applications where interactions within modules or components dominate interactions between modules or components. In addition to the various traditional Markov process instances within computer systems (see, e.g., [8, 10, 9, 6]), these NCD structural properties arise naturally in various traditional aspects of many other domains such as biology, control theory, dynamical systems, economics, genetics, inventory and production systems, queueing theory, and social sciences; refer to, e.g., [5, 22, 11, 4, 9, 6, 3, 30].
@@
where the $\mathbf{P}_{ij}$, $i,j\in[m]$, are nonnegative block matrices in $\mathbb{R}^{n_{i}\times n_{j}}$ such that $\sum_{j\in[m]}\mathbf{P}_{ij}$ are stochastic (i.e., rows sum to one with each element in $[0,1]$), and $\mathbf{P}$ is a stochastic matrix in $\mathbb{R}^{n\times n}$, $n=\sum_{i\in[m]}n_{i}$. Define $\sigma_{i}:=\sum_{\ell=1}^{i-1}n_{\ell}$, $i\in[m]$. Under the assumption that the Markov process is NCD, the elements of the off-diagonal block matrices $\mathbf{P}_{ij}$ are very small relative to the elements of the diagonal block matrices $\mathbf{P}_{ii}$, $i,j\in[m],i\neq j$. In particular, we assume throughout that
$\|\mathbf{P}_{ii}\|=O(1),\qquad\|\mathbf{P}_{ij}\|=O(\epsilon),\qquad\qquad i,j\in[m],\;i\neq j,$ (3)
@@
In addition to their broad applications across a wide range of domains in computing, engineering, science and business, NCD structural properties represent an important fundamental area in the classical theory of Markov processes [22, 27, 28, 18, 19, 7, 14, 24, 30]. Hence, several different (yet related) numerical methods have been developed to efficiently compute the stationary distribution $\bm{\pi}$ of NCD Markov processes, with the most efficient based on different forms of exploiting the aggregation and disaggregation of NCD structural properties as part of an iterative computational method. This collection of the most efficient numerical methods for NCD Markov processes consists predominantly of the algorithms due to Takahashi [27], Vantilborgh [28], and Koury, McAllister and Stewart [18]. On the one hand, although they differ in terms of specific details, these methods exploit aggregation-disaggregation as part of an iterative computational process in a very similar manner and with similar convergence behaviors [7]. On the other hand, despite their significant performance benefits over directly solving (1), these numerical methods can still be prohibitively expensive for computing the stationary distribution of highly large-scale NCD Markov processes and for supporting online or real-time applications of such stochastic processes. Moreover, all of these most efficient methods involve computations in full precision and none of them take advantage of lower precision technology.
@@
Our Contributions. Our focus in this paper is on the design and analysis of numerical methods for computing the stationary distribution $\bm{\pi}$ of NCD Markov processes that address the performance bottlenecks of existing approaches and provide significant computational improvements to support the very large scale of current and emerging computer systems and applications as well as their potential deployment in online or real-time environments. In doing so, we make the following important technical contributions. We first design general classes of computational algorithmic approaches to efficiently determine the solution of (1) for the invariant probability vector $\bm{\pi}$ in (4) of the Markov process $\{X(s)\,;\,s\in\mathbb{Z}_{+}\}$ whose $\mathbf{P}$ matrix has the structural form (2) with properties (3). Each of our general approaches involve a combination of different forms of exploiting advances in computer architectures related to mixed-precision computation, which significantly reduce computation times at the expense of inaccuracies, and exploiting iterative approximate computing methods, which mitigate the impact of inaccurate computations, further reduce computation times, and guarantee convergence. It is important to note that, although low- and mixed-precision computation have received considerable attention in machine learning and related applications, such approaches and methods are far less well established in the context of numerical solutions of stochastic processes. We then derive a mathematical analysis of our general classes of algorithmic approaches that rigorously establishes their important theoretical properties including results on algorithmic approximation errors and convergence behaviors. Lastly, we conduct numerous numerical experiments which empirically demonstrate that our general algorithmic approaches provide orders of magnitude improvements in the computation times to obtain the stationary distribution of NCD Markov processes over the most efficient existing methods which are based on aggregation and disaggregation.
@@
Meanwhile, we note that Horton and Leutenegger [15], motivated by multigrid methods, propose an algorithmic approach for computing the stationary distribution of Markov processes in general without any special structural properties. Although the proposed algorithm is somewhat related to aggregation-disaggregation methods, with the latter methods seen as a special case under specific algorithmic choices, Horton and Leutenegger do not provide any theoretical results on convergence or other algorithmic properties, and they only provide empirical results for Markov processes without NCD structural properties [15]. For these reasons, we do not consider further in this paper the multigrid-based method of Horton and Leutenegger. It is important to note, however, that our general algorithmic approaches devised in this paper within the context of aggregation-disaggregation methods can be analogously exploited within the context of their multigrid-based method for more general Markov processes beyond those having NCD structural properties.
@@
Organization of paper. Section 2 provides technical preliminaries on NCD Markov processes, followed by the presentation of our design of algorithmic solution approaches in Section 3. We derive our mathematical analysis of these algorithmic approaches in Section 4, followed by our proofs of the corresponding theoretical results in Section 5. The empirical results from a representative sample of numerical experiments are presented in Section 6, followed by concluding remarks.
@@ 2 Technical Preliminaries
Beyond our assumptions that the Markov process $\{X(s)\,;\,s\in\mathbb{Z}_{+}\}$ has the structural form (2) with properties (3) and is irreducible and ergodic, we further assume that $z\leq\|\bm{\pi}_{i}\|_{1}$ for a constant $z>0$, and thus from (1) we have $10<z\leq\|\bm{\pi}_{i}\|_{1}<$ 1, $i\in[m]$. The transition probability matrix $\mathbf{P}$, since it is a stochastic matrix, has a dominant unit eigenvalue and corresponding right eigenvector $\mathbf{1}$ given by
$\mathbf{P}\mathbf{1}=\mathbf{1}.$
We additionally assume no other eigenvalues of $\mathbf{P}$ have absolute value one.
@@
i.e., $\mathbf{P}_{ii}$ has only one dominant eigenvalue that is close to one. We further assume that the remaining eigenvalues of $\mathbf{P}_{ii}$ are uniformly bounded away from one. Then, it is well known [19] that there exists a nonsingular matrix $\mathbf{W}_{i}=\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}$ such that
@@
$\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}(\mathbf{I}-\mathbf{P}_{ii})$ $=\begin{pmatrix}{g}_{i}\epsilon+o(\epsilon)&0\\ \mathbf{0}&\mathbf{I}-\mathbf{T}_{i}\end{pmatrix}\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}\qquad\mbox{ and}$
$\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}(\mathbf{I}-\mathbf{P}_{ii})^{-1}$ $=\begin{pmatrix}({g}_{i}\epsilon+o(\epsilon))^{-1}&0\\ \mathbf{0}&(\mathbf{I}-\mathbf{T}_{i})^{-1}\end{pmatrix}\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}.$ (6)
By further imposing the condition
@@
for a constant ${w}_{i}$ independent of $\epsilon$, we guarantee the nondominant eigenvalues $\lambda_{i_{2}},\ldots,\lambda_{i_{n_{i}}}$ of $\mathbf{P}_{ii}$ are uniformly bounded away from one, $i\in[m]$.
@@ 3 Algorithmic Solutions
In this section, we consider various algorithmic approaches for computing the solution of (1) to determine the stationary distribution $\bm{\pi}$ of NCD Markov processes with the goal of providing significant improvements in computational and theoretical properties over the most efficient numerical methods for doing so. Both of our main algorithmic approaches are based on modified variants of the algorithm due to Koury, McAllister and Stewart [18], as the most recent of the three primary aggregation-disaggregation numerical methods, in which we sharply reduce the computational bottlenecks associated with solving systems of linear equations by exploiting various general classes of iterative approximate mixed-precision computing methods. The primary differences between our two main algorithmic approaches concern the aggressiveness of the degree to which we exploit the various general classes of iterative approximate mixed-precision computing methods.
Since both of our algorithmic approaches are based on modified variants of the Koury-McAllister-Stewart (KMS) algorithm, we first present the main steps the original algorithm. The KMS algorithm is composed of an iterative numerical method that starts at iteration 00 with any given initial approximation $\bm{\pi}^{(0)}=(\bm{\pi}_{1}^{(0)},\ldots,\bm{\pi}_{m}^{(0)})$ of the desired solution $\bm{\pi}$. Then, for each subsequent iteration $t=1,2,3,\ldots$, Step $\mathbf{1}$ consists of normalizing the vector components of the current estimate $\bm{\pi}^{(t-1)}$ of the solution $\bm{\pi}$, i.e., compute $\hat{\bm{\pi}}_{i}^{(t-1)}=\bm{\pi}_{i}^{(t-1)}/\|\bm{\pi}_{i}^{(t-1)}\|_{1}$ according to the middle equation in (8), $i\in[m]$. Step $\mathbf{2}$ comprises obtaining the elements of the aggregation matrix $\mathbf{R}^{(t-1)}$ leading up to the current iteration, i.e., compute $\mathbf{R}_{ij}^{(t-1)}=\hat{\bm{\pi}}_{i}^{(t-1)}\mathbf{P}_{ij}\mathbf{1}$ according to the rightmost equation in (8), $i,j\in[m]$. Next, Step $\mathbf{3}$ consists of obtaining the dominant left eigenvector $\mathbf{s}$ of $\mathbf{R}$ by computing the solution of $\mathbf{s}^{(t-1)}=\mathbf{s}^{(t-1)}\mathbf{R}^{(t-1)}$ and $\mathbf{s}^{(t-1)}\mathbf{1}=\|\mathbf{s}^{(t-1)}\|_{1}=1$ according to (9). Step $\mathbf{4}$ comprises computing the Hadamard product $\mathbf{z}^{(t)}=\mathbf{s}^{(t-1)}\odot\hat{\bm{\pi}}^{(t-1)}=(s_{1}^{(t-1)}\hat{\bm{\pi}}_{1}^{(t-1)},\ldots,s_{m}^{(t-1)}\hat{\bm{\pi}}_{m}^{(t-1)})$. Then, Step $\mathbf{5}$ consists of solving the sequence of $m$ systems of linear equations $\bm{\pi}_{i}^{(t)}=\bm{\pi}_{i}^{(t)}\mathbf{P}_{ii}+\sum_{j<i}\mathbf{z}_{j}^{(t)}\mathbf{P}_{ji}+\sum_{j>i}\bm{\pi}_{j}^{(t)}\mathbf{P}_{ji}$ in the order $i=m,\ldots,1$, thus rendering the blockwise estimates of the components of $\bm{\pi}^{(t)}$ for iteration $t$. Step $\mathbf{6}$ comprises conducting a test for convergence, and halting with the return of the current estimate $\bm{\pi}^{(t)}$ if the solution is sufficiently accurate; otherwise, the iteration index $t$ is incremented and the iterative process of Step $\mathbf{1}$ – Step $\mathbf{6}$ is repeated.
We now present each of our two main algorithmic approaches in turn.
3.1 Conservatively Modified Koury-McAllister-Stewart Algorithm
The performance bottlenecks of the KMS algorithm concern computing the solution of $m+1$ systems of linear equations in each iteration, specifically the sequence of $n_{i}\times n_{i}$ linear systems in Step $\mathbf{5}$, $i\in[m]$, as well the $m\times m$ linear system in Step $\mathbf{3}$ for problems with large $m$. To address these performance bottlenecks, our first general class of modified variants of the KMS algorithm consists of replacing the standard linear solvers in Step $\mathbf{5}$ and Step $\mathbf{3}$ with general classes of approximate mixed-precision computing methods, each of which involves a combination of different forms of lower precision technology to reduce computation costs and different forms of iterative approximate methods to further reduce computation costs with adjustments to realize full-precision results. Instances of the former include forms of multi-precision arithmetic or stochastic rounding, whereas instances of the latter include forms of iterative refinement [29], Richardson iteration [13], and generalized minimal residual (GMRES) [21] methods.

### v2

@@ Abstract
In this paper we consider the problem of computing the stationary distribution of nearly completely decomposable Markov processes, a well-established area in the classical theory of Markov processes with broad applications in the design, modeling, analysis and optimization of computer systems and applications. We devise a general mathematical framework of numerical solution methods that exploits forms of mixed-precision computation to significantly reduce computation times and that exploits forms of iterative approximate computing approaches to mitigate the impact of inaccurate computations, further reduce computation times, and ensure convergence. Then we derive a mathematical analysis that establishes theoretical properties of our general algorithmic framework including results on approximation errors, convergence behaviors, and other algorithmic characteristics. Numerical experiments demonstrate that our general algorithmic framework provides significant improvements in computation times over the most-efficient existing numerical methods.
keywords:
nearly completely decomposable Markov processes , mathematical analysis , mixed-precision computation , iterative approximate methods , numerical analysis
[…]
organization=Mathematics of Computation, IBM Research,addressline=Thomas J. Watson Research Center, city=Yorktown Heights, postcode=10598, state=NY, country=USA
@@ 1 Introduction
Markov processes often play an important role in the design and mathematical modeling, analysis and optimization of a wide range of computer systems and applications. A critically important goal in the mathematical modeling, analysis and optimization of these computer systems and applications as well as their dynamic behavior concerns the need to determine the stationary distribution of the corresponding Markov process, from which various measures and quantities of interest can in turn be directly obtained to facilitate and support such design, analysis and optimization of these computer systems and applications.
More precisely, consider a discrete-time Markov process $\{X(s)\,;\,s\in\mathbb{Z}_{+}\}$ defined on the state space $[n]:=\{1,\ldots,n\}$, and let ${\mathbf{P}}\in\mathbb{R}^{n\times n}$ and $\bm{\pi}:=(\pi_{1},\ldots,\pi_{n})$ respectively denote its transition probability matrix and stationary distribution. Here, $\mathbb{R}$ and $\mathbb{Z}_{+}$ denote the set of real numbers and nonnegative integers, respectively. Assuming this Markov process is irreducible and ergodic, then the invariant probability vector $\bm{\pi}$ exists and is uniquely determined as the solution of [1, 2]
@@
One particularly important class of structural properties often arising in such Markov processes concerns the so-called “aggregation of variables” [3] in which all the states (variables) of the system can be grouped into a relatively small number of relatively large clusters such that the magnitude of transition probabilities between states within a cluster overwhelmingly dominates the relatively tiny magnitude of transition probabilities between states in different clusters. Markov processes with these structural properties, called nearly completely decomposable (NCD), make it possible to effectively investigate the system and its dynamics by exploiting the analysis of interactions among the states within each individual cluster without reference to the interactions among the states across clusters and by exploiting the analysis of interactions among the states across different clusters without reference to the interactions among the states within clusters. Such NCD properties arise naturally in many computer systems and applications due to the prevalence of levels of modular abstraction or hierarchies of components in the design of these computer systems and applications, where interactions within modules or components dominate interactions between modules or components. In addition to the various traditional Markov process instances within computer systems and applications (see, e.g., [4, 5, 6, 7]), these NCD structural properties arise naturally in various traditional aspects of many other domains such as queueing theory, control theory, dynamical systems, inventory and production systems, economics, biology, genetics, and social sciences; refer to, e.g., [8, 3, 9, 10, 6, 7, 11, 12].
@@
where $\mathbf{P}_{ij}$, $i,j\in[m]$, are nonnegative block matrices in $\mathbb{R}^{n_{i}\times n_{j}}$ such that $\sum_{j\in[m]}\mathbf{P}_{ij}$ are stochastic (i.e., rows sum to one with each element in $[0,1]$), and $\mathbf{P}$ is a stochastic matrix in $\mathbb{R}^{n\times n}$, $n=\sum_{i\in[m]}n_{i}$. Define $\sigma_{i}:=\sum_{\ell=1}^{i-1}n_{\ell}$, $i\in[m]$. Under the assumption that the Markov process is NCD, the elements of the off-diagonal block matrices $\mathbf{P}_{ij}$ are very small relative to the elements of the diagonal block matrices $\mathbf{P}_{ii}$, $i,j\in[m],i\neq j$. In particular, we assume throughout that
$\|\mathbf{P}_{ii}\|=O(1)\qquad\mbox{ and }\qquad\|\mathbf{P}_{ij}\|=O(\epsilon),\qquad\qquad i,j\in[m],\;i\neq j,$ (3)
@@
In addition to their broad applications across a wide range of domains in computing, engineering, science and business, NCD structural properties represent an important fundamental area in the classical theory of Markov processes [3, 13, 14, 15, 16, 17, 18, 19, 12]. Hence, several different (yet related) numerical methods have been developed to efficiently compute the stationary distribution $\bm{\pi}$ of NCD Markov processes, with the most efficient based on different forms of exploiting the aggregation and disaggregation of NCD structural properties as part of an iterative computational method. This collection of the most-efficient numerical methods for NCD Markov processes consists predominantly of the algorithms due to Takahashi [13], Vantilborgh [14], and Koury, McAllister and Stewart [15]. On the one hand, although they differ in terms of specific details, all these methods exploit aggregation-disaggregation as part of an iterative computational process in a very similar manner and with similar convergence behaviors [17]. On the other hand, despite their significant performance benefits over directly solving (1), these numerical methods can still be prohibitively expensive for computing the stationary distribution of highly large-scale NCD Markov processes and for supporting online or real-time applications of such stochastic processes. Moreover, all of these most-efficient methods involve computations in full precision and none of them take advantage of advances in mixed-precision technology.
@@
Our Contributions. Our focus in this paper is on the design and analysis of a general mathematical framework of numerical methods for computing the stationary distribution $\bm{\pi}$ of NCD Markov processes that addresses the performance bottlenecks of existing algorithms and provides significant computational improvements to support the very large scale of current and emerging computer systems and applications as well as their potential deployment in online or real-time environments. In doing so, we make the following important technical contributions. We first devise a general algorithmic framework of computational approaches to efficiently determine the solution of (1) for the invariant probability vector $\bm{\pi}$ in (4) of the Markov process $\{X(s)\,;\,s\in\mathbb{Z}_{+}\}$ whose $\mathbf{P}$ matrix has the structural form (2) with properties (3). Our general mathematical framework involves a combination of different forms of exploiting advances in computer architectures related to mixed-precision computation, which significantly reduce computation times at the expense of inaccuracies, and exploiting advances in iterative approximate computing methods, which mitigate the impact of inaccurate computations, further reduce computation times, and guarantee convergence. It is important to note that, although mixed-precision computation has received considerable attention in machine learning and related applications, such approaches and methods are far less well established in the context of numerical solutions of stochastic processes; moreover, our framework combines such mixed-precision computation with iterative approximate computing methods. We then derive a mathematical analysis of our general algorithmic framework that rigorously establishes their important theoretical properties including results on approximation errors, convergence behaviors, and other algorithmic characteristics. Lastly, we conduct many numerical experiments which empirically demonstrate that our general algorithmic framework provides orders of magnitude improvements in the computation times to obtain the stationary distribution of NCD Markov processes over the most-efficient existing methods which exploit aggregation-disaggregation. These numerical experiments are based on simulation of different graphics processing unit (GPU) architectures that we validated against physical experiments on a particular GPU architecture.
@@
Meanwhile, we note that Horton and Leutenegger [20], motivated by multigrid methods, propose an algorithmic approach for computing the stationary distribution of Markov processes in general without any special structural properties. Although the proposed algorithm is somewhat related to aggregation-disaggregation algorithms, with the latter methods seen as a special case under specific algorithmic choices, Horton and Leutenegger do not provide any theoretical results on convergence or other algorithmic properties, and they only provide empirical results for Markov processes without NCD structural properties [20]. For these reasons, we do not consider further in this paper the multigrid-based algorithm of Horton and Leutenegger. It is important to note, however, that the numerical methods of our general mathematical framework devised in this paper within the context of aggregation-disaggregation algorithms can be analogously exploited within the context of their multigrid-based algorithm for more general Markov processes beyond those having NCD structural properties.
@@
Organization of paper. Section 2 provides technical preliminaries on NCD Markov processes, followed by the presentation of our design of a general mathematical solution framework together with a few of its representative examples in Section 3. We derive our mathematical analysis of these examples of our algorithmic framework in Section 4, followed by our proofs of the corresponding theoretical results in Section 5. The empirical results from a representative sample of our numerical experiments are presented in Section 6, followed by concluding remarks.
@@ 2 Technical Preliminaries
Beyond our assumptions that the Markov process $\{X(s)\,;\,s\in\mathbb{Z}_{+}\}$ has the structural form (2) with properties (3) and is irreducible and ergodic, we further assume that $\tau\leq\|\bm{\pi}_{i}\|_{1}$ for some constant $\tau>0$, and thus from (1) we have $10<\tau\leq\|\bm{\pi}_{i}\|_{1}<$ 1, $i\in[m]$. This assumption of $\tau\leq\|\bm{\pi}_{i}\|_{1}$ simply ensures that none of the $\bm{\pi}_{i}$ go to 00 as $\epsilon$ goes to 00, $i\in[m]$, which is a standard regularity condition [19]. The transition probability matrix $\mathbf{P}$, since it is a stochastic matrix, has a dominant unit eigenvalue and corresponding right eigenvector $\mathbf{1}$ given by $\mathbf{P}\mathbf{1}=\mathbf{1}$. We additionally assume no other eigenvalues of $\mathbf{P}$ have absolute value one.
@@
i.e., the dominant eigenvalue of $\mathbf{P}_{ii}$ is close to one. Let us further assume that the remaining eigenvalues of $\mathbf{P}_{ii}$ are uniformly bounded away from one, whose underlying condition guaranteeing this assumption will be justified below. Then, it is well known [16] that there exists a nonsingular matrix $\mathbf{W}_{i}=\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}$ such that
@@
$\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}(\mathbf{I}-\mathbf{P}_{ii})=\begin{pmatrix}{g}_{i}\epsilon+o(\epsilon)&0\\ \mathbf{0}&\mathbf{I}-\mathbf{T}_{i}\end{pmatrix}\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}\qquad\mbox{ and }\qquad\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}(\mathbf{I}-\mathbf{P}_{ii})^{-1}=\begin{pmatrix}({g}_{i}\epsilon+o(\epsilon))^{-1}&0\\ \mathbf{0}&(\mathbf{I}-\mathbf{T}_{i})^{-1}\end{pmatrix}\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}.$ (6)
By imposing the condition
@@
for a constant ${w}_{i}$ independent of $\epsilon$, we guarantee that the nondominant eigenvalues $\lambda_{i_{2}},\ldots,\lambda_{i_{n_{i}}}$ of $\mathbf{P}_{ii}$ are uniformly bounded away from one, $i\in[m]$, as assumed above.
@@ 3 Algorithmic Solutions
In this section, we first devise a general mathematical framework of numerical methods for computing the solution of (1) to determine the stationary distribution $\bm{\pi}$ of NCD Markov processes with the goal of providing significant improvements in computational and theoretical properties over the most-efficient numerical methods for doing so. While our general algorithmic framework of computational approaches can be applied within the context of any of these existing numerical methods and beyond, we then consider representative examples of applying our algorithmic framework within the context of the algorithm due to Koury, McAllister and Stewart [15] (KMS), since it is the most recent of the most-efficient numerical methods. In particular, we consider two such examples of our general algorithmic framework in which we sharply reduce the computational bottlenecks associated with solving systems of linear equations by exploiting general classes of iterative approximate mixed-precision computing methods. The primary differences between these two representative examples of our algorithmic framework concern the degree of aggressiveness with which we exploit the general classes of iterative approximate mixed-precision computing methods.
3.1 General Mathematical Framework
All the computational approaches to compute the stationary distribution of NCD Markov processes, including the most-efficient aggregation-disaggregation algorithms such as KMS, involve numerical methods to solve systems of linear equations, generically denoted here by $\mathbf{A}\mathbf{x}=\mathbf{b}$, where the numerical solution of such linear systems represents the core computational bottlenecks of these approaches. Our general mathematical framework consists of a combination of universal forms of exploiting: (1) advances in computer architectures and technology related to mixed-precision computation, which significantly reduce computation times at the expense of inaccurate results; and (2) advances in iterative approximate computing methods, which mitigate the impact of inaccurate computations, further reduce computation times, and guarantee convergence. More precisely, each individual invocation of standard full-precision linear system solvers to compute occurrences of $\mathbf{A}\mathbf{x}=\mathbf{b}$ is replaced by a sequence of computational steps based on general classes of iterative approximate mixed-precision methods for numerically solving linear systems. This involves each computation of $\mathbf{A}\mathbf{x}=\mathbf{b}$ in a chosen mixed-precision mode supported by advanced computer architectures and technology as part of a selected iterative approximate numerical computing method. The former chosen mixed-precision mode significantly reduces computation times at the cost of inaccurate results, instances of which include forms of multi-precision arithmetic or stochastic rounding; whereas the latter selected iterative approximation numerical method addresses the inaccurate computations of mixed-precision modes, further reduces computation times and ensures convergence, instances of which include forms of iterative refinement (IR) [21], Richardson iteration (RI) [22], and generalized minimal residual (GMRES) [23] methods.
Since both representative examples of applying our general algorithmic framework herein are conducted within the context of the KMS algorithm, we first present the main steps the original algorithm. In particular, the KMS algorithm is composed of an iterative numerical method that starts at iteration 00 with any given initial approximation $\bm{\pi}^{(0)}=(\bm{\pi}_{1}^{(0)},\ldots,\bm{\pi}_{m}^{(0)})$ of the desired solution $\bm{\pi}$. Then, for each subsequent iteration $t=1,2,3,\ldots$:
Step $\mathbf{1}$ consists of normalizing the vector components of the current estimate $\bm{\pi}^{(t-1)}$ of the solution $\bm{\pi}$, i.e., compute $\hat{\bm{\pi}}_{i}^{(t-1)}=\bm{\pi}_{i}^{(t-1)}/\|\bm{\pi}_{i}^{(t-1)}\|_{1}$ according to the middle equation in (8), $i\in[m]$.
Step $\mathbf{2}$ comprises obtaining the elements of the aggregation matrix $\mathbf{R}^{(t-1)}$ leading up to the current iteration, i.e., compute $\mathbf{R}_{ij}^{(t-1)}=\hat{\bm{\pi}}_{i}^{(t-1)}\mathbf{P}_{ij}\mathbf{1}$ according to the rightmost equation in (8), $i,j\in[m]$.
Step $\mathbf{3}$ consists of obtaining the dominant left eigenvector $\mathbf{s}$ of $\mathbf{R}$ by computing the solution of $\mathbf{s}^{(t-1)}=\mathbf{s}^{(t-1)}\mathbf{R}^{(t-1)}$ and $\mathbf{s}^{(t-1)}\mathbf{1}=\|\mathbf{s}^{(t-1)}\|_{1}=1$ according to (9).
Step $\mathbf{4}$ comprises computing the Hadamard product $\mathbf{z}^{(t)}=\mathbf{s}^{(t-1)}\odot\hat{\bm{\pi}}^{(t-1)}=(s_{1}^{(t-1)}\hat{\bm{\pi}}_{1}^{(t-1)},\ldots,s_{m}^{(t-1)}\hat{\bm{\pi}}_{m}^{(t-1)})$.
Step $\mathbf{5}$ consists of solving the series of $m$ systems of linear equations $\bm{\pi}_{i}^{(t)}=\bm{\pi}_{i}^{(t)}\mathbf{P}_{ii}+\sum_{j<i}\mathbf{z}_{j}^{(t)}\mathbf{P}_{ji}+\sum_{j>i}\bm{\pi}_{j}^{(t)}\mathbf{P}_{ji}$ in the order $i=m,\ldots,1$, thus rendering the blockwise estimates of the components of $\bm{\pi}^{(t)}$ for iteration $t$.
Step $\mathbf{6}$ comprises a test for convergence, halting with the return of the current estimate $\bm{\pi}^{(t)}$ if the solution is sufficiently accurate; otherwise, the iteration index $t$ is incremented and the iterative process of Step $\mathbf{1}$ – Step $\mathbf{6}$ is repeated.
The core performance bottlenecks of the KMS algorithm concern computing the solution of $m+1$ systems of linear equations in each iteration $t$, specifically the series of $n_{i}\times n_{i}$ linear systems in Step $\mathbf{5}$, $i\in[m]$, as well the $m\times m$ linear system in Step $\mathbf{3}$ for problems with large $m$. In strong contrast, Step $\mathbf{1}$ and Step $\mathbf{2}$ simply set up the computation in Step $\mathbf{3}$ and involve relatively little computation compared to the solution of such a linear system in Step $\mathbf{3}$; similarly, Step $\mathbf{4}$ simply sets up the computation in Step $\mathbf{5}$ and involves relatively little computation compared to the solution of such linear systems in Step $\mathbf{5}$. We now present in turn two representative examples of applying our general algorithmic framework to address the core performance bottlenecks of Step $\mathbf{5}$ and Step $\mathbf{3}$.
3.2 Representative Conservative Example of General Algorithmic Framework
The first representative example of a general application of our algorithmic framework within the context of the KMS algorithm consists of replacing each invocation of a full-precision linear system solver in Step $\mathbf{5}$ and Step $\mathbf{3}$ with a combination of the selection of an appropriate level of mixed-precision supported by the available computer architecture and the selection of any appropriate instance of the general classes of iterative approximate numerical methods that jointly realize full-precision results. This representative example of our general algorithmic framework is conservative in the sense that the iterative approximate mixed-precision computing method solves each occurrence of a linear system through a sequence of iterations which lead to full-precision accuracy (though more efficiently). We summarize such a representative conservative example in Algorithm 1 where the IR method is chosen for the iterative approximate numerical computing method of our general mathematical framework.